OpenStudy (anonymous):

combining like terms simplify each expression -4+7(1-3m) = -7m+8

OpenStudy (anonymous):

-4 + 7(1 - 3m) = -4 + 7 + 7 * -3m = 3 - 21m

OpenStudy (anonymous):

Jimmyrep the -4 and the -3 are like terms correct?

OpenStudy (anonymous):

right

OpenStudy (lilg132):

oops i just solved it sorry :p

OpenStudy (anonymous):

which you add -4+-3 correct

OpenStudy (anonymous):

following on from last post 3 - 21m = 7m + 8 - right i was starting to solve it too

OpenStudy (anonymous):

lol

OpenStudy (lilg132):

-4 + 7 - 21m = -7m + 8 3 - 21m = 7m + 8

OpenStudy (anonymous):

ok thank you lol

OpenStudy (anonymous):

correction ifaison - 4 and 7 are like terms - -4 and -3m are not like terms

OpenStudy (anonymous):

Jimmyrep lilg132 dont confuse me I am a rookie lol trying to get ready for an asessment test